For a 1GHz processor computer, compute clock cycle time.
Answer #1

Clock cycle time is defined as the reciprocal of clock frequency.

In your case, the clock frequency is 1GHz which is 109 Hz.

Hence, clock cycle time = 1/frequency = 1/109 = 10-9 = 1.0 ns(nano seconds).